OpenStudy (anonymous):

Find the equation of the parabola whose axis is parallel to the y-axis, with vertex (1,3) and containing the point (5,7)

OpenStudy (anonymous):

y = a(x-1)^2 + 3 7 = a (5-1)^2 + 3 7 = 16a + 3 a = 1/4 y = 1/4 (x-1)^2 + 3

OpenStudy (anonymous):

Thank you!